40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Sat Feb 20 19:11:22 2016, Koji, Electronics, Characterization, Dark current measurement of the HQE PD and other PDs P2197992.jpgPD_dark_current.pdfPD_dark_current_others.pdf
    Reply  Sun Mar 13 21:22:27 2016, Koji, Electronics, Characterization, Dark current measurement of the HQE PD and other PDs 
Message ID: 251     Entry time: Sat Feb 20 19:11:22 2016     Reply to this: 253
Author: Koji 
Type: Electronics 
Category: Characterization 
Subject: Dark current measurement of the HQE PD and other PDs 

Dark current of the HQE PD and other PDs were measured.

- The HQE PDs were loaded on the new PD transportation cages (Attachment 1)
The PDs are always shorted by a clean PD plugs. The PD element is still capped with Kapton seals.

- The assignment of the container/slot and the PDs are as follows

Slot \ Container A B C D E
1 A1-23 B1-22 C1-07 C1-11 C1-17
2 A1-25 B1-23 C1-08 C1-12 C1-21
3 B1-01 C1-03 C1-09 C1-14 D1-08
4 B1-16 C1-05 C1-10 C1-15 D1-10

- The measurement has been done with KEITHLEY sourcemeter SMU2450.

- The result is shown in Attachment 2. Most of the PDs show the dark current of ~3nA at 15V bias. C1-05 and C1-07 showed higher dark current at high V region. We should avoid using them for the aLIGO purpose. I hope they are still OK at low bias V if there is no noise issue (TBC). You can not read the PD names on the plot for the nominal ones, but that's OK as they are almost equivalent.

- As a comparison, the dark current of a C30655 (serial #10) was measured. Considering a DC current due to an anbient light (although the PD was covered), the dark current of the HQE PD seems double of C30655.

- Taking an advantage of having the setup, I took the same measurement for the Laser Comp. PDs in ATF. I gave the identification as #1 and #2. #1 has full-length legs while #2 has trancated legs. As Zach reported before, they showed significantly high dark current. (Attachment 3)

Attachment 1: P2197992.jpg  2.567 MB  | Hide | Hide all | Show all
P2197992.jpg
Attachment 2: PD_dark_current.pdf  25 kB  Uploaded Fri Mar 11 19:56:26 2016  | Hide | Hide all | Show all
PD_dark_current.pdf
Attachment 3: PD_dark_current_others.pdf  14 kB  Uploaded Fri Mar 11 19:56:37 2016  | Show | Hide all | Show all
ELOG V3.1.3-